OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)) METAL WILL BE REWARDED

OpenStudy (anonymous):

g(2)=(2)^2-6(2)-7=-15 f(-15)=-15+8=-7